OpenStudy (anonymous):

I need the equation of a line that has a slope of -3 and goes through the points (-4,5) in standard form help?

OpenStudy (anonymous):

first set up in point-slope (y-y1)=m(x-x1) y-5=-3(x+4) distribute y-5=-3x-12 add 5 to both sides y=-3x-7 add 3x to both sides 3x+y=-7